# Generic derivations, differential largeness, and NTP2

### v1

@@ 1. Introduction
In this note, $\mathcal{L}$ is a language extending $\mathcal{L}_{\operatorname{ring}}=\{0,1,+,\cdot\}$, and $K$ is an $\mathcal{L}$-structure expanding a field of characteristic zero. We let $n,m,k,r$ range over $\mathbb{N}=\{0,1,2,\ldots\}$. Let $\dim$ denote algebraic dimension on subsets of cartesian powers of $K$, so $\dim X$ is the dimension of the Zariski closure of $X\subseteq K^{n}$.
The last few years have seen new approaches to the model theory of “tame” differential fields, often under the guiding principle: “The model theory of a differential field is completely determined by the model theory of the underlying field, so long as the derivation is generic.” Genericity roughly means that anything that can happen, subject to the constraints imposed by differential algebra and the theory of the underlying field, does happen.
@@
Many tame classes of fields are known to be algebraically bounded, including real closed fields, algebraically closed fields, and henselian fields of characteristic zero. If $K$ is algebraically bounded, then $\dim$ is a definable dimension on $K$, meaning that for any definable family $(X_{a})_{a\in A}$ and any $d$, the set of $a\in A$ for which $\dim(X_{a})=d$ is definable. Algebraic boundedness was introduced by van den Dries [23], and the equivalence of van den Dries’ definition with the one given here is [13, Lemma 2.12].
@@
When $K$ is algebraically bounded, genericity is first-order axiomatizable. Fornasiero and Terzo go on to show that if $K$ is algebraically bounded and $\delta$ is generic, then several model-theoretic properties of $K$ transfer to $(K,\delta)$, including quantifier elimination and model completeness.
@@
It seems difficult to answer this question without some understanding of the definable sets in $K$. In Proposition 2.3 below, we give a “topological” axiomatization of differential largeness in terms of the étale open topology of Johnson, Tran, Walsberg, and Ye [12]. In [24], Walsberg and Ye isolated the class of éz-fields—large, algebraically bounded fields that are “topologically tame” with respect to this topology. For these fields, we have a good enough understanding of definable sets to answer the question positively.
@@ 3. Transferring NTP2 and NATP
Assume that $T^{\delta}_{g}$ has TP2, as witnessed by an $\mathcal{L}^{\delta}$-formula $\varphi(x,y)$ and an array of tuples $(a_{i,j})_{i,j<\omega}$. By [4, Lemma 3.2], we may also assume that $|x|=1$. By Fact 1.4(2), the formula $\varphi$ is equivalent to a formula of the form $\psi(\nabla^{r}x,\nabla^{s}y)$ for natural numbers $r,s$ and an $\mathcal{L}$-formula $\psi$. By replacing $a_{i,j}$ by $\nabla^{s}(a_{i,j})$ and augmenting $y$, we may assume that $s=1$. For the rest of this proof, we fix $r$ and an $\mathcal{L}$-formula $\psi(x_{0},\ldots,x_{r},y)$ (each $x_{i}$ unary) such that $\psi(\nabla^{r}x,y)$ has TP2. We assume that $r$ is minimal with this property. Note that if $r=0$, then $\psi$ is an $\mathcal{L}$-formula, so $T$ has TP2 and we are done. Thus, we assume for the remainder of the proof that $r>0$. We fix an array $(a_{i,j})_{i,j<\omega}$ witnessing TP2, and we may arrange that this array is strongly $\mathcal{L}^{\delta}$-indiscernible, meaning that each row $(a_{i,j})_{j<\omega}$ is $\mathcal{L}^{\delta}$-indiscernible over the other rows and that the sequence of rows is $\mathcal{L}^{\delta}$-indiscernible.
@@
Then $A_{\varnothing}=A$ and for $B\subseteq 2^{<\omega}$, the set $\bigcup_{\eta\in B}A_{\eta}$ is an antichain if and only if $B$ is an antichain. Set
@@
For $n\in\mathbb{N}$, we set $n\langle 0\rangle\coloneqq\langle 0,0,\ldots,0\rangle\in 2^{n}$ (so $0\langle 0\rangle=\varnothing$). We set $F_{n}\coloneqq\pi\big{(}\bigcap_{i\leqslant n}X_{i\langle 0\rangle}\big{)}$, we let $Z_{n}$ denote the Zariski closure of $F_{n}$, and we take $m$ with $Z_{m}=Z_{n}$ for $n\geqslant m$. Arguing as in the proof of Theorem 3.1, we have that $m>0$, that $\dim(Z_{m})<r$, and that
@@
for all $m\langle 0\rangle\lhd\eta_{0}\lhd\eta_{1}\lhd\cdots\lhd\eta_{m}$ (this uses that $(b_{\eta})_{\eta\in C}$ and $(b_{\nu})_{\nu\in C^{\prime}}$ have the same $\mathcal{L}^{\delta}$-type for any two finite chains $C,C^{\prime}\subseteq 2^{<\omega}$ of the same length, as a consequence of strong indiscernibility). Now for each $\nu\in 2^{<\omega}$, let $c_{\nu}\coloneqq(a_{\varnothing},\ldots,a_{m\langle 0\rangle},a_{m\langle 0\rangle^{\frown}\nu})$ and set
@@ Acknowledgments
Part of this research were conducted while E. Kaplan was hosted by the Max Planck Institute for Mathematics, and he thanks the MPIM for its support and hospitality. This research was supported in part by the National Science Foundation under Award No. DMS-2103240.

### v2

@@ 1. Introduction
In this note, $\mathcal{L}$ is a language extending $\mathcal{L}_{\operatorname{ring}}=\{0,1,+,\cdot\}$, and $K$ is an $\mathcal{L}$-structure expanding a field of characteristic zero. We let $n,m,k,r$ range over $\mathbb{N}=\{0,1,2,\ldots\}$.
The last few years have seen new approaches to the model theory of “tame” differential fields, often under the guiding principle: “The model theory of a differential field is largely determined by the model theory of the underlying field, so long as the derivation is generic.” Genericity roughly means that any system of differential equations and inequalities has a solution, so long as the existence of a solution is consistent with the theory of the underlying field.
@@
Many tame classes of fields are known to be algebraically bounded, including real closed fields, algebraically closed fields, and henselian fields of characteristic zero (in the language of rings, though one can also allow a predicate for the valuation ring in the last case). Algebraic boundedness was introduced by van den Dries [23], and the equivalence of van den Dries’ definition with the one given here is [13, Lemma 2.12].
Let $\dim$ denote algebraic dimension on subsets of cartesian powers of $K$, so $\dim X$ is the transcendence degree of the function field $K(X)$ over $K$ for $X\subseteq K^{n}$. Equivalently, $\dim X=\max_{a\in X^{*}}\operatorname{trdeg}(a|K)$, where $(K^{*},X^{*})$ is a $|K|^{+}$-saturated elementary extension of $(K,X)$; see [23, Section 2]. If $K$ is algebraically bounded, then $\dim$ is a definable dimension on $K$, meaning that for any definable family $(X_{a})_{a\in A}$ and any $r$, the set of $a\in A$ for which $\dim(X_{a})=r$ is definable.
@@
When $K$ is algebraically bounded, genericity is first-order axiomatizable. Fornasiero and Terzo go on to show that if $K$ is algebraically bounded and $\delta$ is generic, then several model-theoretic properties of $K$ transfer to $(K,\delta)$, including quantifier elimination and model completeness [8, Corollaries 3.16 and 3.17].
@@
It seems difficult to answer this question without some understanding of the definable sets in $K$. In Proposition 2.3 below, we give a “topological” axiomatization of differential largeness in terms of the étale open topology of Johnson, Tran, Walsberg, and Ye [12]. Widawski gives a similar axiomatization of differential largeness in his thesis [25] (see Remark 2.6), and we thank the referee for bringing this to our attention.
In [24], Walsberg and Ye isolated the class of éz-fields—large, algebraically bounded fields that are “topologically tame” with respect to the étale open topology. For these fields, we have a good enough understanding of definable sets to answer the question positively.
@@ 3. Transferring NTP2 and NATP
Assume that $T^{\delta}_{g}$ has TP2, as witnessed by an $\mathcal{L}^{\delta}$-formula $\varphi(x,y)$ and an array of tuples $(a_{i,j})_{i,j<\omega}$. By [4, Lemma 3.2], we may also assume that $|x|=1$. By Fact 1.4(2), the formula $\varphi$ is equivalent to a formula of the form $\psi(\nabla^{r}x,\nabla^{s}y)$ for natural numbers $r,s$ and an $\mathcal{L}$-formula $\psi$. By replacing $a_{i,j}$ by $\nabla^{s}(a_{i,j})$ and augmenting $y$, we may assume that $s=0$. For the rest of this proof, we fix $r$ and an $\mathcal{L}$-formula $\psi(x_{0},\ldots,x_{r},y)$ (each $x_{i}$ unary) such that $\psi(\nabla^{r}x,y)$ has TP2. We assume that $r$ is minimal with this property. Note that if $r=0$, then $\psi$ is an $\mathcal{L}$-formula, so $T$ has TP2 and we are done. Thus, we assume for the remainder of the proof that $r>0$. We fix an array $(a_{i,j})_{i,j<\omega}$ witnessing TP2, and we may arrange that this array is strongly $\mathcal{L}^{\delta}$-indiscernible, meaning that each row $(a_{i,j})_{j<\omega}$ is $\mathcal{L}^{\delta}$-indiscernible over the other rows and that the sequence of rows is $\mathcal{L}^{\delta}$-indiscernible; see [15, Lemma 5.6].
@@
Then $A_{\emptyset}=A$ and for $B\subseteq 2^{<\omega}$, the set $\bigcup_{\eta\in B}A_{\eta}$ is an antichain if and only if $B$ is an antichain. Set
@@
For $n\in\mathbb{N}$, we set $n\langle 0\rangle\coloneqq\langle 0,0,\ldots,0\rangle\in 2^{n}$ (so $0\langle 0\rangle=\emptyset$). We set $F_{n}\coloneqq\pi\big(\bigcap_{i\leqslant n}X_{i\langle 0\rangle}\big)$, we let $Z_{n}$ denote the Zariski closure of $F_{n}$, and we take $m$ with $Z_{m}=Z_{n}$ for $n\geqslant m$. Arguing as in the proof of Theorem 3.1, we have that $m>0$, that $\dim(Z_{m})<r$, and that
@@
for all $m\langle 0\rangle\lhd\eta_{0}\lhd\eta_{1}\lhd\cdots\lhd\eta_{m}$ (this uses that $(b_{\eta})_{\eta\in C}$ and $(b_{\nu})_{\nu\in C^{\prime}}$ have the same $\mathcal{L}^{\delta}$-type for any two finite chains $C,C^{\prime}\subseteq 2^{<\omega}$ of the same length, as a consequence of strong indiscernibility). Now for each $\nu\in 2^{<\omega}$, let $c_{\nu}\coloneqq(a_{\emptyset},\ldots,a_{m\langle 0\rangle},a_{m\langle 0\rangle^{\frown}\nu})$ and set
@@ Acknowledgments
The first author is a postdoctoral researcher of the Fonds de la Recherche Scientifique – FNRS, and was supported in part by the National Science Foundation under Award No. DMS-2103240. Parts of this research were conducted while the first author was hosted by the Max Planck Institute for Mathematics, and he thanks the MPIM for its support and hospitality. We thank the referee for helpful feedback.